A taxicab charges $4.25 for the first mile and $1.50 for each addtional mile. Write a rule for describing the total rate r as a function of the total miles m. What is the taxi rate for 12 miles? My answer is: 4.25 + 1.50m, $20.75
Answer by Maths68(1474)

A taxicab charges $4.25 for the first mile and $1.50 for each addtional mile. Write a rule for describing the total rate r as a function of the total miles m. What is the taxi rate for 12 miles? My answer is: 4.25 + 1.50m, $20.75
Charges for first 1 mile = 4.25
Charges for each addtional mile = 1.5
Taxi Rate = r
Total Miles = m

Taxi Rate = Charges for First mile + 1.5 (Total Miles - 1)
r=4.25+1.5(m-1)
r=4.25+1.5m-1.5
r=2.75+1.5m (Formula for calculating such rate)

For 12 Miles
=============
Charges for first 1 mile = 4.25
Charges for each addtional mile = 1.5
Taxi Rate = r
Total Miles = m = 12

r=2.75+1.5m
r=2.75+1.5(12)
r=2.75+18
r=20.75
Charges for 12 miles ride $20.75